OpenStudy (anonymous):

if a polygon have two interior angle of 90 degrees and the others are each 120 degrees then what is the number of sides of the polygon?

OpenStudy (mathmate):

Fact: The exterior angles of a polygon add up to 360 degrees. An exterior angle is 180-interior angle in degrees. Let the polygon have n sides, then 2(180-90) + (n-2)(180-120) = 360 60(n-2)=360-180=180 n-2=3 n=5 So it is a pentagon